OpenStudy (anonymous):

119=n-66 please solve

OpenStudy (anonymous):

Add 66 to both sides. n = 185

OpenStudy (anonymous):

n - 66 = 119 +66 +66 n = 185

OpenStudy (anonymous):

and 5p=140 ?

OpenStudy (anonymous):

Lol just divide 140 by 5